OpenStudy (anonymous):

solve the linear system using the addition-elimination method 3x-5y=-2 -2x+4y=1

OpenStudy (cwrw238):

multiply the first equation by 2 and second one by 3 adding these 2 new equations will eliminate x then solve for y finally find x by plugging value of y into first equation

OpenStudy (anonymous):

i cant figure out the values, every time i try i get different numbers that dont work in the equation

OpenStudy (cwrw238):

ok ill start it for you 6x - 10y = -4 thats first equation x 2 - remember to multiply each term by 2 now second by 3:- -6x + 12y = 3 adding 2y = - 1

OpenStudy (cwrw238):

do u follow that?

OpenStudy (anonymous):

yeah i get that its just when i try to find x and plug both numbers back to check if its right it doesnt work out

OpenStudy (cwrw238):

y = 1/2 right?

OpenStudy (anonymous):

i got y= -1/2

OpenStudy (anonymous):

and x = 1 1/2 which doesnt work in the equation

OpenStudy (cwrw238):

good - u spotted my deliberate mistake now you plug y = -1/2 into 3x - 5y = -2

OpenStudy (cwrw238):

3x - 5(-1/2) = -2 3x = -2 - 5/ 2 right?

OpenStudy (anonymous):

its 3x = -4 1/2 and x = (-4 1/2)/3

OpenStudy (cwrw238):

3x = -9/2 x = -9/2 * 1/3 = -9/6 = -3/2

OpenStudy (cwrw238):

yea

OpenStudy (cwrw238):

-3/2 = -1 1/2

OpenStudy (anonymous):

right, but when i plug that in it doesnt work. it comes to -3.5 + 2.5 = -2

OpenStudy (anonymous):

oh i just realized what i did wrong. thank for your help :-)